Two Kleaning ladies company needs a 50% bleach mixture solution. they make 12 liter at a time from a 40% bleach solution and a 70% bleach solution. How many liters do they need of each?
Answer by lwsshak3(11628)

Two Kleaning ladies company needs a 50% bleach mixture solution. they make 12 liter at a time from a 40% bleach solution and a 70% bleach solution. How many liters do they need of each?
**
let x=liters of 40% solution to be used for the mixture
12-x=liters of 70% solution to be used for the mixture
..
40%x+70%(12-x)=50%*12
.4x+8.4-.7x=6
.3x=2.4
x=8
12-x=4
ans:
liters of 40% solution to be used for the mixture=4
liters of 70% solution to be used for the mixture=8